OpenStudy (anonymous):

what is dy/dx: x^2=x-y/x+y

OpenStudy (anonymous):

Implicit D!

OpenStudy (anonymous):

\[2x = \frac{(x-y)^{\prime}(x+y)-(x+y)^{\prime}(x-y)}{(x+y)^2}\] \[2x = \frac{(1-y^{\prime})(x+y)-(1+y^{\prime})(x-y)}{(x+y)^2}\] Now just collect everything and factor out the y prime

OpenStudy (anonymous):

how do I factor out the y prime?